OpenStudy (anonymous):

The radius of a circle is 2.9 miles. What is the circle's area? Use 3.14 for π and round your answer to the nearest hundredth.

OpenStudy (anonymous):

please help me! :)))))))))))))))

OpenStudy (diyadiya):

Area Of a Circle = Pi * r ^2

OpenStudy (diyadiya):

3.14 * (2.9)^2 =3.14 * 8.41 =26.4

OpenStudy (anonymous):

3.14* 2.9*2.9 = 26.4074 rounded would be 26.41

OpenStudy (anonymous):

ok thanks guys

OpenStudy (anonymous):

so would it be 26.41 or 26.4?

OpenStudy (anonymous):

if you want to round to the nearest hundredth, it would be 26.41

OpenStudy (anonymous):

ok thanks

OpenStudy (diyadiya):

26.41

OpenStudy (lilg132):

26.4207941865

OpenStudy (lilg132):

nearest hundreth wouldn't it be 26.42?